OpenStudy (anonymous):

Find a polynomial f(x) satisfying the equation xf''(x) + 3f(x) = 4x^2.

OpenStudy (anonymous):

Since f''(x) and f'(x) has the maximum power (order) 2. Hence we can say the ploynomial is of 3rd order. Let the polynomial be \[f(x) = ax ^{3}+bx ^{2}+cx+d\] This gives, \[f'(x) = 3ax ^{2}+2bx+c\] And differentiating once more, \[f''(x) = 6ax+2b\] Hence, \[xf''(x)+3f'(x) = 6ax^{2}+2bx+9ax ^{2}+6bx+3c = 12ax ^{2}+8bx+3c\] From the differential equation \[xf''(x)+3f'(x) = 12ax ^{2}+8bx+3c = 4x ^{2} \] => \[12a = 4, 8b=0, 3c = 0 => a = \frac{1}{3}, b=0, c=0 \] Hence the polynomial is \[f(x) = \frac{1}{3}x ^{2}+d\] Where, d is any constant or constant of integration

OpenStudy (anonymous):

im so lost, my homework said it was wrong, any thing else that could be the answer?

OpenStudy (anonymous):

Is it \(xf''(x) + 3f(x) = 4x^2?\) or \(xf''(x) + 3f'(x) = 4x^2?\)

OpenStudy (anonymous):

just f(x)

OpenStudy (anonymous):

Find a polynomial f(x) that satisfies xf ''(x) + 3f (x) = 4x^2.

OpenStudy (anonymous):

Oh I see.

OpenStudy (anonymous):

yeah this is calculus... and im lostulus lol

OpenStudy (anonymous):

I wasn't focusing. It's doable.

OpenStudy (anonymous):

i sure hope so

OpenStudy (anonymous):

The answer is \(f(x)=\frac{4}{3}x^2-\frac{8}{9}x\).

OpenStudy (anonymous):

wow... however did you do it?

OpenStudy (anonymous):

Here how it's done. We're going to assume that the solution is a quadratic polynomial, since we have the sum of the function and its derivative adds up to \(4x^2\). So let \(f(x)=ax^2+bx+c \implies f'(x)=2ax+b \text{ and } f''(x)=2a.\) Plug all these into the given DE to get \(2ax+3ax^2+3bx+3c=4x^2\), or \(3ax^2+(2a+3b)x+3c=4x^2\). From the above equation the left hand side should have a coffienet of \(4\) for \(x^2\), and all other coffienents must be \(0\). That is \(3a=4\), \(2a+3b=0\) and \(3c=0\). You can easily solve the system above and find \(a=\frac{4}{3}, b=-\frac{8}{9}\) and \(c=0\). Thus \(f(x)=\frac{4}{3}x^2-\frac{8}{9}x\).

OpenStudy (anonymous):

Don't hisetate to ask if there's anything you don't get.

OpenStudy (anonymous):

okay so how did you get 2ax +3ax^2 +3bx + 3c = 4x^2

OpenStudy (anonymous):

thanks for youre help, but I have to go somewhere, you can still explain if you would like too, and i promise you ill come back and read it. thanks again!

OpenStudy (anonymous):

We are given that \[xf''(x) + 3f(x) = 4x^2\] But based on our assumption \(f(x)=ax^2+bx+c\) and \(f''(x)=2a\). Substitute for f(x) and f''(x) in the equation to get that.

OpenStudy (anonymous):

@AnwarA The f(x) is a cubic polynomial